Mikebroome:

Annette’s paddle boat travels 2 km/h in still water.the boat is paddled 4 km downstream in the same time that it takes to go 1 km upstream. What is the speed of the river

Mercury:

the paddle boat travels 2 km/hr in still water. let's call the speed of the river r. upstream, the boat goes 2 - r km/hr downstream, the boat goes 2 + r km/hr now, we can use d = st (distance = speed*time) to solve for r. upstream: 1 = (2 - r)t downstream: 4 = (2 + r)t t is the same for both, because both journeys take the same time. therefore, you can solve both equations for t, set them equal to each other, and solve for r.